Analyze at what degree to cut an apple to share with your friends, if your friends ate that apple at what velocity would the apple slice arrive at its destination. As the apple is arriving at its destination traveling at the speed of light how many hours would it take for a neko to travel to their destination of home, please convert your answer to milliseconds .
1
Expert's answer
2020-04-05T13:08:10-0400

Imagine that we have N=36 friends and the apple have a shape of an ideal sphere. Thus, we would have to cut the slices 10 degrees each:


"\theta_\text{slice}=\frac{360^\circ}{N}=\frac{360^\circ}{36}=10^\circ."


Now, imagine that we have 36 friends 24 light years away from the Earth, but due to technical limitations all we can is to ship one slice every 24 hours. Therefore, one slice will reach a friend in


"t_1=\frac{d}{c}=\frac{24\text{ years}\cdot c}{c}=24 \text{ years}."


In 24 years, the first slice only will reach the first friend. The second slice will reach the second friend 24 hours later, the third - 48 hours (2 days) after the first slice, the fourth - 3 days later, therefore, the last of the 36 slices will reach the 36th friend in


"\Delta t=N-1=36-1=35 \text{ days}."


The last slice will reach the last recipient 35 days later than the first slice. Therefore, the total time is


"t=t_1+\Delta t=\text{24 years 35 days},"


or


"t=24\cdot365.22\cdot24\cdot3600+35\cdot24\cdot3600=\\=760344192\text{ s}."